OpenStudy (anonymous):

solve the exponential equation 9^2x+1=11 (2x+1) is all raised

OpenStudy (anonymous):

\[9^{2x+1} = 11 \Rightarrow \log_9(9^{2x+1}) = \log_9(11) \Rightarrow (2x+1)\log_99 = \log_911\]

OpenStudy (anonymous):

\[2x+1 = \log_911 \Rightarrow 2x = \log_9(11) - 1 \Rightarrow x = \frac{1}{2}(\log_9(11)-1)\]

OpenStudy (anonymous):

we need the value for x

OpenStudy (anonymous):

its not going to be pretty, its not like x = 5, or 8. You would have to plug that into a calculator.

OpenStudy (anonymous):

or wolframalpha.com

OpenStudy (anonymous):

you should be able to calculate value of x from that equation. spoon feeding is bad for teaching/learning any subject.

OpenStudy (anonymous):

2xln9=ln11 2x=ln(2) x=ln(2)^1/2